OpenStudy (anonymous):

Factor completely: 4x2 + 25x + 6 (4x + 1)(x + 6) (4x + 6)(x + 1) (2x + 3)(2x + 2) (2x + 6)(2x + 1)

OpenStudy (anonymous):

The first one.

OpenStudy (anonymous):

sure??

OpenStudy (anonymous):

yep.

OpenStudy (anonymous):

Want the steps?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

4x^(2)+25x+6 For a polynomial of the form ax^(2)+bx+c, find two factors of a*c (24) that add up to b (25).In this problem 6*(1)/(4)=(6)/(4) (which is (c)/(a)) and 6+(1)/(4)=(25)/(4) (which is ((b)/(a)) , so insert 6 as the right hand term of one factor and (1)/(4) as the right-hand term of the other factor. (x+6)(x+(1/4)) Remove the fraction by multiplying the first term of the factor by the denominator of the second term. (x+6)(4x+1)

OpenStudy (anonymous):

ok thank you